# Machine Learning-Based Design and Monte Carlo Simulation of a Neutron Beam Shutter for Cyclotron-Based Neutron Sources

### v1

@@ Abstract
This study proposes a novel design methodology for neutron beam shutters that integrates Monte Carlo simulations (MCNP) with machine learning techniques to enhance shielding performance and accelerate the design process. The target facility is a compact neutron science platform where neutrons are produced by proton beams from a cyclotron striking a neutron production target. The system includes both thermal and fast neutron beamlines. A beam shutter is installed on the thermal neutron line to reduce occupational radiation exposure during maintenance activities.
@@
Note: This work has been submitted to Nuclear Engineering and Technology for peer review.
@@ 5 Conclusion
CRediT authorship contribution statement
Li-Fang Chen: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Neutron beam shutter simulation, Data curation, Writing – original draft.
Jyh-Shen Tsay,Cheng-Si Tsao: Writing – review and editing.
Sheng-Long Jeng: Supervision, Project administration, Funding acquisition.

### v3

# Machine Learning-Assisted Optimization of Modular Neutron Shielding Based on Monte Carlo Simulations

@@ Abstract
This study proposes a novel design methodology for neutron beam shutters that integrates Monte Carlo simulations (MCNP) with machine learning techniques to enhance shielding performance and accelerate the design process. The target facility is a compact neutron science platform where neutrons are produced by proton beams from a cyclotron striking a neutron production target. A beam shutter is installed on the thermal neutron line to reduce occupational radiation exposure during maintenance activities.
@@ 4 Discussion
4.1 Layer-wise Material Composition and Performance Interpretation
Figure 9 presents the statistical analysis of material composition across all 25 layers for the top and bottom 5% of shielding configurations, ranked by total neutron flux predicted from 1,000 randomly generated designs using the trained machine learning model. Material indices shown in the figure correspond to the standardized classification scheme adopted from Reference [12].
The analysis reveals that high-performance shielding configurations—those with the lowest predicted neutron flux—exhibit a strong preference for iron in the front layers (i.e., those closest to the neutron source), effectively moderating fast neutrons. Polyethylene and borated polyethylene are periodically distributed in the middle and rear layers, providing neutron moderation and absorption. In contrast, low-performance configurations (with the highest flux values) tend to contain high proportions of polyethylene and borated polyethylene across all layers, lacking sufficient high-density materials in the initial layers to attenuate fast neutrons.
These results suggest that although polyethylene-based materials offer favorable moderation and absorption properties, their low density limits their standalone effectiveness in comprehensive shielding. Strategic layering—introducing high-density materials in the early stages, followed by appropriately arranged moderator and absorber layers—is essential for achieving substantial reductions in neutron flux and enhancing overall shielding performance.